Find the length s of an arc that subtends a central angle of 311° in a circle of radius 2 m.
s =
Please explain how this is solved
Thank you

Answer by Alan3354(69443)

Find the length s of an arc that subtends a central angle of 311° in a circle of radius 2 m.
---------------
Circumference = 2*pi*r
C = 2*pi*r
----
s = C*(311/360)